# Hairy black holes in extended Einstein-Maxwell-scalar theories with magnetic charge and kinetic couplings

### v1

@@ Abstract
We study black hole (BH) solutions in extended Einstein-Maxwell-scalar theories which is classified in a subclass of the $U(1)$ gauge-invariant scalar-vector-tensor theories. The scalar field is coupled to the vector field, which has electric and magnetic charges. For the static and spherically symmetric spacetime, we investigate modifications to the Reissner-Nordström solutions focusing on the three types of scalar-vector interactions, including derivative couplings. We solve the field equations analytically in two asymptotic regions which are the vicinity of the BH horizon and the spatial infinity, and clarify the condition for the existence of scalar hair. To understand the behaviors of solutions in intermediate scales, the field equations are integrated numerically for concrete models with different types of couplings. We find new hairy BH solutions with scalar hair in the presence of magnetic charge and kinetic coupling. The magnetic charge plays an important role in distinguishing hairy BH solutions originated from three types of different interactions at a large coupling limit.
@@ I Introduction
In recent years, the observations of the nonlinear regime of gravity have represented a significant development. For example, the direct detection of gravitational waves from compact binary coalescences by LIGO-Virgo-KAGRA Collaborations Abbott et al. (2016a, b, 2017a, 2017b, 2017c, 2017d, 2019, 2020a, 2020b, 2020c, 2021, 2023) and the imaging of black hole (BH) shadow by the Event Horizon Telescope Akiyama et al. (2019, 2022) opened up a new window for exploring physics around BHs. In the Einstein-Maxwell system without matter, asymptotically flat and stationary BH solutions are described by only three parameters, i.e., mass, electric charge, and angular momentum Israel (1967); Carter (1971); Ruffini and Wheeler (1971); Hawking (1972). Hence, the Kerr-Newman BH Newman et al. (1965) is the unique solution in such a system. This statement is known as the “no-hair theorem.” It is natural to wonder whether this theorem still holds even in the presence of a new degree of freedom. If a canonical scalar field minimally coupled to gravity is introduced to the Einstein-Maxwell system, then, the profile of scalar field settles down to be trivial Chase (1970); Bekenstein (1972, 1996) , and the spacetime is still described by the Kerr-Newman solutions. This fact shows the absence of a “scalar hair” due to the validity of the no-hair theorem even in the presence of the canonical scalar field. Constructing BH solutions with the “scalar hair” in the context of GR is not a simple problem (see Ref. Herdeiro and Radu (2015) for review). However, if a scalar field couples to the Maxwell field, it can take a nontrivial profile Gibbons (1982); Gibbons and Maeda (1988); Garfinkle et al. (1991); Campbell et al. (1991); Lee and Weinberg (1991); Reuter (1992); Monni and Cadoni (1996); Stefanov et al. (2007a, b, 2008); Sheykhi and Hajkhalili (2014); Sheykhi and Kazemi (2014); Sheykhi et al. (2015); Fan and Lu (2015); Dehghani (2018); Herdeiro et al. (2018); Myung and Zou (2019a); Herdeiro and Oliveira (2019); Myung and Zou (2019b); Konoplya and Zhidenko (2019); Hod (2019); Dehghani and Setare (2019); Filippini and Tasinato (2019); Boskovic et al. (2019); Fernandes et al. (2019); Astefanesei et al. (2019); Herdeiro and Oliveira (2020); Fernandes (2020); Brihaye et al. (2020); Promsiri et al. (2023). Such a coupling between scalar and vector fields arises in several theories, e.g., effective Lagrangians of axion Peccei and Quinn (1977); Georgi et al. (1986) which is a strong candidate of dark matter and compactified higher-dimensional theories Gibbons (1982); Gibbons and Maeda (1988); Garfinkle et al. (1991).
@@
The interactions between scalar and vector fields can be generalized requiring the equations of motion being up to second order in terms of derivatives for the absence of Ostrogradsky instabilities Ostrogradsky (1850); Woodard (2015). Indeed, the application of this method to a single scalar degree of freedom nonminimally coupled to gravity leads the Horndeski theory, which is the most general scalar-tensor theory with second-order equations of motion Horndeski (1974); Deffayet et al. (2011); Kobayashi et al. (2011). For the massive vector degree of freedom, the same approach results in the generalized Proca theory Heisenberg (2014); Tasinato (2014a, b). The Horndeski and the generalized Proca theories can be unified as the scalar-vector-tensor (SVT) theories in association with new interactions between the scalar field and the vector field Heisenberg (2018). If the vector field satisfies the $U(1)$ gauge symmetry especially, it reduces to the so-called $U(1)$ gauge-invariant SVT theories in which the Einstein-Maxwell-scalar theory is classified as a subclass.
@@
The interactions between scalar field and vector field in the $U(1)$ gauge-invariant SVT theories are described by three Lagrangians denoted as $\mathcal{L}_{2},\mathcal{L}_{3}$, and $\mathcal{L}_{4}$. The scalar or axionic-type couplings to the vector field discussed in Refs. Gibbons and Maeda (1988); Garfinkle et al. (1991); Campbell et al. (1991); Lee and Weinberg (1991); Reuter (1992); Sheykhi and Hajkhalili (2014); Fan and Lu (2015); Myung and Zou (2019a); Boskovic et al. (2019); Myung and Zou (2019b); Dehghani and Setare (2019); Filippini and Tasinato (2019); Fernandes et al. (2019); Astefanesei et al. (2019); Herdeiro and Oliveira (2020); Fernandes (2020) correspond to particular cases of the $\mathcal{L}_{2}$ interaction111 The existence of hairy BH solutions arising from the interactions $\mathcal{L}_{3}$ and $\mathcal{L}_{4}$ are discovered in Ref. Heisenberg and Tsujikawa (2018), and the stability of these solutions against odd-parity perturbations are studied in Ref. Heisenberg et al. (2018).. In these works, the scalar field $\phi$ couples to the vector field through the scalar type interaction of the form $g_{1}(\phi)F_{\mu\nu}F^{\mu\nu}$ or the axionic-type interaction of the form $g_{2}(\phi)F_{\mu\nu}\tilde{F}^{\mu\nu}$ where $F_{\mu\nu}$ is the strength tensor of vector field and $\tilde{F}_{\mu\nu}$ is its dual, $g_{1}$ and $g_{2}$ are functions of $\phi$. On the other hand, in the range of the $U(1)$ gauge-invariant SVT theories, one can consider other types of interactions. We can take account of the field derivatives $\nabla_{\alpha}\phi$, instead of the field $\phi$ itself, coupled to the vector field as $\nabla_{\alpha}\phi\nabla^{\alpha}\phi\,F_{\mu\nu}F^{\mu\nu}$ and $\nabla_{\mu}\phi\nabla^{\nu}\phi\,F^{\mu\alpha}F_{\nu\alpha}$ where $\nabla_{\mu}$ is the covariant derivative operator. If we demand the $U(1)$ gauge field to be static and spherically symmetric at the time of configuration, however, the latter interaction, $\nabla_{\mu}\phi\nabla^{\nu}\phi\,F^{\mu\alpha}F_{\nu\alpha}$, reduces to be in proportion to the former interaction, $\nabla_{\alpha}\phi\nabla^{\alpha}\phi\,F_{\mu\nu}F^{\mu\nu}$. This shows that the two types of interactions are degenerated on the configuration and cannot be distinguished. Moreover, we cannot examine the effect of a scalar product $F_{\mu\nu}\tilde{F}^{\mu\nu}$ at the background since it identically vanishes.
@@ II $U(1)$ gauge-invariant SVT theories and field equations
where $\phi$, $V$ are functions of the radial coordinate $r$, and $P$ corresponds to the magnetic charge. From Eqs. (2.4) and (2.5), the scalar quantities $X$, $F$, $\tilde{F},$ and $Y$, defined in Eqs. (2.2)-(2.3), are evaluated as
@@
which belongs to the subclass of theories given in Eq. (2.1). Here, $g_{1},g_{2}$, and $\bar{g}_{3}$ are arbitrary functions of the scalar field $\phi$ and its kinetic term $X$. For the later convenience, we normalize $\bar{g}_{3}(\phi,X)$ as
@@
without loss of generality.
@@ III.1 Solutions around the horizon
where $f_{i}$, $h_{i}$, $V_{0}$, $V_{i}$, $\phi_{0}$, $\phi_{i}$ are constants. Hereafter, the constant $V_{0}$ will be omitted by setting $V_{0}=0$ without loss of generality, since $V$ always appears with a derivative in the field equations by virtue of the $U(1)$ gauge symmetry. We treat the effect of the couplings $g_{1},g_{2},g_{3}$ as corrections to the RN metric with a magnetic charge $P$ given by
@@
$\hskip 20.0pt\left.-{3\Phi_{1,\phi X}\over 8\left(\Phi_{1,X}-2r_{h}^{4}\right)}+{\Phi_{2,\phi}\over 2\gamma}\left[{\Phi_{2,\phi}\over\Phi_{1,\phi}}-{3\Phi_{2,X}\over 2(\Phi_{1,X}-2r_{h}^{4})}\right]\right)\phi_{1}^{2}+\left({(1-\mu)\left(\gamma_{g_{1}g_{3}}\Phi_{1,XX}+2\Phi_{2,X}^{2}\right)\over 4r_{h}\gamma_{g_{1}g_{3}}\left(\Phi_{1,X}-2r_{h}^{4}\right)}\right)\phi_{1}^{3}\,,$ (3.8)
@@
The coupling functions $g_{1}$, $g_{2}$, $g_{3}$ are evaluated on the horizon with $\left.(\phi,X)\right|_{r=r_{h}}=(\phi_{0},0)$. These couplings give rise to corrections to the metric components at the order of $(r-r_{h})^{2}$. When the coupling does not depend on the field $\phi$ itself, the coefficient $\phi_{1}$ in Eq. (3.4) identically vanishes which results in the elimination of $\phi_{2}$ given in Eq. (3.8). In such a case, we can numerically confirm that the scalar field settles down to be constant $\phi=\phi_{0}$ in all the region outside the horizon which implies the absence of scalar hair. As a consequence, we find that the $\phi$-dependence in the functions $g_{i}(\phi,X)$ is essential for achieving hairy BH solutions.
@@ IV Concrete models and numerical analyses
For the scalar and the axionic-type couplings of the forms $g_{1}F$ and $g_{2}\tilde{F}$, respectively, it has shown that the charged BH solutions with scalar hair exist Gibbons and Maeda (1988); Garfinkle et al. (1991); Lee and Weinberg (1991); Boskovic et al. (2019); Astefanesei et al. (2019); Fernandes et al. (2019). This subsection aims to examine whether the new BH solutions arise from the derivative interaction, $\bar{g}_{3}Y$, in intermediate scales. For this purpose, we concentrate on the case $c_{3}=1$ with vanishing $c_{2}$, $c_{3}$, and $d_{i}$. We solve the field equations given in Eqs. (2.9)-(2.12) numerically with the model parameters $\mu=0.1$ and $P=0.1M_{\rm pl}r_{h}$. The analytic solutions around the horizon given in Eqs. (3.3)-(3.8) are used as boundary conditions with $\phi_{0}=-1.0\times 10^{-4}M_{\rm pl}$ at $r=1.001r_{h}$. We present the result in Fig. 1.
Figure 1: Numerical solutions of $f,h,V,\phi,h-f$ outside the horizon for the coupling $c_{3}=1$, with vanishing $c_{2}$, $c_{3}$, and $d_{i}$. The boundary conditions are chosen to satisfy Eqs. (3.3)-(3.8) with $\mu=0.1$, $P=0.1M_{\rm pl}r_{h}$, $\phi_{0}=-1.0\times 10^{-4}M_{\rm pl}$ at $r=1.001r_{h}$.

### v2

@@ Abstract
We study static and spherically symmetric black hole (BH) solutions in extended Einstein-Maxwell-scalar theories, which are classified in a subclass of the $U(1)$ gauge-invariant scalar-vector-tensor theories. The scalar field is coupled to the vector field, which has electric and magnetic charges. We investigate modifications to the Reissner-Nordström solution focusing on the three types of scalar-vector interactions, including derivative couplings. We solve the field equations analytically in two asymptotic regions which are the vicinity of the BH horizon and the spatial infinity, and clarify the condition for the existence of scalar hair. To understand the behaviors of solutions in intermediate scales, the field equations are integrated numerically for concrete models with different types of couplings. We find new hairy BH solutions with scalar hair in the presence of magnetic charge and kinetic coupling. The magnetic charge plays an important role in distinguishing hairy BH solutions originated from three types of different interactions at a large coupling limit.
@@ I Introduction
Experimental tests of General Relativity (GR) is well performed in the solar system Will (2014). As a result, it is known that the weak gravitational field can be described by GR in high precision. On the other hand, the accuracy of GR in the strong field regime has been not yet clarified well enough. In recent years, the observations of the nonlinear regime of gravity have represented a significant development. For example, the direct detection of gravitational waves from compact binary coalescences by LIGO-Virgo-KAGRA Collaborations Abbott et al. (2016a, b, 2017a, 2017b, 2017c, 2017d, 2019, 2020a, 2020b, 2020c, 2021, 2023) and the imaging of black hole (BH) shadow by the Event Horizon Telescope Akiyama et al. (2019, 2022) opened up a new window for exploring physics around BHs. The action including Maxwell electromagnetic field besides GR is called Einstein-Maxwell theory. The authors of Ref. Colléaux et al. (2024) have classified the generalized higher-order Einstein-Maxwell Lagrangians, which consist of terms linear in the Riemann tensor and quadratic in derivatives of the field strength tensor. In the four dimensional Einstein-Maxwell system without matter, asymptotically flat and stationary BH solutions are described by only three parameters, i.e., mass, electric charge, and angular momentum Israel (1967); Carter (1971); Ruffini and Wheeler (1971); Hawking (1972). Hence, the Kerr-Newman BH Newman et al. (1965) is the unique solution in such a system. This statement is known as the “no-hair theorem.” It is natural to wonder whether this theorem still holds even in the presence of a new degree of freedom. If a canonical scalar field minimally coupled to gravity is introduced to the Einstein-Maxwell system, then, the profile of scalar field settles down to be trivial Chase (1970); Bekenstein (1972, 1996) , and the spacetime is still described by the Kerr-Newman solutions. This fact shows the absence of a “scalar hair” due to the validity of the no-hair theorem even in the presence of the canonical scalar field. Constructing BH solutions with the “scalar hair” in the context of GR is not a simple problem Chew et al. (2023); Chew and Lim (2024); Ghosh et al. (2023); Chew and Yeom (2024) (see Ref. Herdeiro and Radu (2015) for review). However, if a scalar field couples to the Maxwell field, it can take a nontrivial profile Gibbons (1982); Gibbons and Maeda (1988); Garfinkle et al. (1991); Campbell et al. (1991); Lee and Weinberg (1991); Reuter (1992); Monni and Cadoni (1996); Stefanov et al. (2007a, b, 2008); Sheykhi and Hajkhalili (2014); Sheykhi and Kazemi (2014); Sheykhi et al. (2015); Fan and Lu (2015); Dehghani (2018); Herdeiro et al. (2018); Myung and Zou (2019a); Herdeiro and Oliveira (2019); Myung and Zou (2019b); Konoplya and Zhidenko (2019); Hod (2019); Dehghani and Setare (2019); Filippini and Tasinato (2019); Boskovic et al. (2019); Fernandes et al. (2019); Astefanesei et al. (2019); Herdeiro and Oliveira (2020); Brihaye et al. (2020); Fernandes (2020); Mahapatra et al. (2020); Priyadarshinee et al. (2021); Priyadarshinee and Mahapatra (2023); Promsiri et al. (2023); Belkhadria and Pombo (2023). Such a coupling between scalar and vector fields arises in several theories, e.g., effective Lagrangians of axion Peccei and Quinn (1977); Georgi et al. (1986) which is a strong candidate of dark matter and compactified higher-dimensional theories Gibbons (1982); Gibbons and Maeda (1988); Garfinkle et al. (1991).
@@
The interactions between scalar and vector fields can be generalized requiring the equations of motion being up to second order in terms of derivatives for the absence of Ostrogradsky instabilities Ostrogradsky (1850); Woodard (2015). Indeed, the application of this method to a single scalar degree of freedom nonminimally coupled to gravity leads the Horndeski theory, which is the most general scalar-tensor theory with second-order equations of motion Horndeski (1974); Deffayet et al. (2011); Kobayashi et al. (2011)111 Horndeski theories can be further generalized so as to allow the existence of higher order derivatives while keeping the number of propagating degree of freedom by virtue of degeneracy conditions. This class of theories is called Degenerate Higher-Order Scalar-Tensor (DHOST) theories Langlois and Noui (2016).. For the massive vector degree of freedom, the same approach results in the generalized Proca theory Heisenberg (2014); Tasinato (2014a, b). The Horndeski and the generalized Proca theories can be unified as the scalar-vector-tensor (SVT) theories in association with new interactions between the scalar field and the vector field Heisenberg (2018). If the vector field satisfies the $U(1)$ gauge symmetry especially, it reduces to the so-called $U(1)$ gauge-invariant SVT theories in which the Einstein-Maxwell-scalar theory is classified as a subclass.
@@
The interactions between scalar field and vector field in the $U(1)$ gauge-invariant SVT theories are described by three Lagrangians denoted as $\mathcal{L}_{2},\mathcal{L}_{3}$, and $\mathcal{L}_{4}$. The scalar or axionic-type couplings to the vector field discussed in Refs. Gibbons and Maeda (1988); Garfinkle et al. (1991); Campbell et al. (1991); Lee and Weinberg (1991); Reuter (1992); Sheykhi and Hajkhalili (2014); Fan and Lu (2015); Myung and Zou (2019a); Boskovic et al. (2019); Myung and Zou (2019b); Dehghani and Setare (2019); Filippini and Tasinato (2019); Fernandes et al. (2019); Astefanesei et al. (2019); Herdeiro and Oliveira (2020); Fernandes (2020) correspond to particular cases of the $\mathcal{L}_{2}$ interaction222 The existence of hairy BH solutions arising from the interactions $\mathcal{L}_{3}$ and $\mathcal{L}_{4}$ are discovered in Ref. Heisenberg and Tsujikawa (2018), and the stability of these solutions against odd-parity perturbations are studied in Ref. Heisenberg et al. (2018).. In these works, the scalar field $\phi$ couples to the vector field through the scalar type interaction of the form $g_{1}(\phi)F_{\mu\nu}F^{\mu\nu}$ or the axionic-type interaction of the form $g_{2}(\phi)F_{\mu\nu}\tilde{F}^{\mu\nu}$ where $F_{\mu\nu}$ is the strength tensor of vector field and $\tilde{F}_{\mu\nu}$ is its dual, $g_{1}$ and $g_{2}$ are functions of $\phi$. On the other hand, in the range of the $U(1)$ gauge-invariant SVT theories, one can consider other types of interactions. We can take account of the field derivatives $\nabla_{\alpha}\phi$, instead of the field $\phi$ itself, coupled to the vector field as $\nabla_{\alpha}\phi\nabla^{\alpha}\phi\,F_{\mu\nu}F^{\mu\nu}$ and $\nabla_{\mu}\phi\nabla^{\nu}\phi\,F^{\mu\alpha}F_{\nu\alpha}$ where $\nabla_{\mu}$ is the covariant derivative operator. If we demand the $U(1)$ gauge field to be static and spherically symmetric at the time of configuration, however, the latter interaction, $\nabla_{\mu}\phi\nabla^{\nu}\phi\,F^{\mu\alpha}F_{\nu\alpha}$, reduces to be in proportion to the former interaction, $\nabla_{\alpha}\phi\nabla^{\alpha}\phi\,F_{\mu\nu}F^{\mu\nu}$. This shows that the two types of interactions are degenerated on the configuration and cannot be distinguished. Moreover, we cannot examine the effect of a scalar product $F_{\mu\nu}\tilde{F}^{\mu\nu}$ at the background since it identically vanishes.
@@ II $U(1)$ gauge-invariant SVT theories and field equations
where $\phi$, $V$ are functions of the radial coordinate $r$, and $P$ corresponds to the monopolar magnetic charge. From Eqs. (2.4) and (2.5), the scalar quantities $X$, $F$, $\tilde{F},$ and $Y$, defined in Eqs. (2.2)-(2.3), are evaluated as
@@
which belongs to the subclass of theories given in Eq. (2.1). Here, $g_{1},g_{2}$, and $\bar{g}_{3}$ are arbitrary functions of the scalar field $\phi$ and its kinetic term $X$. In the following sections, we will search for nontrivial BH solutions such that $f$, $h$, $\phi$, and $V$ are regular at the horizon. Assuming that the above Lagrangian is also regular at the horizon, $g_{1}$ and $g_{2}$ need to be functions that have zero or positive powers of $X$ since this quantity vanishes at the horizon (see Eq. (2.6)). On the other hand, the situation is different for $g_{3}$. As can be seen from Eq. (2.6), the quantity $Y$ is proportional to $X$ in the form $Y=(2hV^{\prime 2}/f)X$. This shows that, even for $\bar{g}_{3}\propto X^{-1}$, the combination $\bar{g}_{3}Y$ is regular at the horizon. In order to make this point clear, we normalize $\bar{g}_{3}$ as
@@
without loss of generality. If we demand $g_{3}$ has zero or positive power of $X$, the Lagrangian $\sqrt{-g}\,g_{3}Y/(4X)$ can be regular at the horizon.
@@ III.1 Solutions around the horizon
where $f_{i}$, $h_{i}$, $V_{0}$, $V_{i}$, $\phi_{0}$, $\phi_{i}$ are constants. Hereafter, the constant $V_{0}$ will be omitted by setting $V_{0}=0$ without loss of generality, since $V$ always appears with a derivative in the field equations by virtue of the $U(1)$ gauge symmetry. Similar to the regularity in Eq. (3.1), we assume that the coupling functions are regular at the horizon. Since the kinetic term $X$ is expanded around the horizon as $X=-(h_{1}\phi_{1}^{2}/2)(r-r_{h})+{\cal O}((r-r_{h})^{2})$, the expansion of coupling $g_{i}(\phi,X)$ with $i=1,2,3$ can be expressed as $g_{i}(\phi,X)=g_{i}(\phi_{0},0)+[g_{i,\phi}(\phi_{0},0)-g_{i,X}(\phi_{0},0)h_{1}\phi_{1}^{2}/2](r-r_{h})+{\cal O}((r-r_{h})^{2})$. We denote $g_{i}(\phi_{0},0)$, $g_{i,\phi}(\phi_{0},0)$, $g_{i,X}(\phi_{0},0)$ as $g_{i}$, $g_{i,\phi}$, $g_{i,X}$ for simplicity in the following discussion.
We treat the effect of the couplings as corrections to the RN metric with a magnetic charge $P$ given by
@@
$\hskip 20.0pt\left.-{3\Phi_{1,\phi X}\over 8\left(\Phi_{1,X}-2r_{h}^{4}\right)}+{\Phi_{2,\phi}\over 2\gamma_{g_{1}g_{3}}}\left[{\Phi_{2,\phi}\over\Phi_{1,\phi}}-{3\Phi_{2,X}\over 2(\Phi_{1,X}-2r_{h}^{4})}\right]\right)\phi_{1}^{2}+\left({(1-\mu)\left(\gamma_{g_{1}g_{3}}\Phi_{1,XX}+2\Phi_{2,X}^{2}\right)\over 4r_{h}\gamma_{g_{1}g_{3}}\left(\Phi_{1,X}-2r_{h}^{4}\right)}\right)\phi_{1}^{3}\,,$ (3.8)
@@
The coupling functions $g_{1}$, $g_{2}$, $g_{3}$ are evaluated on the horizon with $(\phi,X)|_{r=r_{h}}=(\phi_{0},0)$. We note that the authors of Ref. Babichev et al. (2017) have obtained hairy BH solutions with non-zero $\phi$ and $X$ at the horizon in the shift symmetric beyond Horndeski theories. In such theories, by virtue of the shift symmetry, the vanishing Noether current gives rise to the solutions with $\phi^{\prime}$ divergent at the horizon while the kinetic term $X$ remains finite there. In the scope of this paper, we focus on the theories described by Eq. (2.7) in which the shift symmetry of scalar field does not hold due to the existence of $\phi$-dependence in the coupling functions $g_{1}$, $g_{2}$, and $g_{3}$. Hence, we focus on the solutions endowed with scalar field, which is regular at the horizon. As seen in Eq. (3.5), the couplings give rise to corrections to the metric components at the order of $(r-r_{h})^{2}$. When the coupling does not depend on the field $\phi$ itself, the coefficient $\phi_{1}$ in Eq. (3.4) identically vanishes which results in the elimination of $\phi_{2}$ given in Eq. (3.8). In such a case, we can numerically confirm that the scalar field settles down to be constant $\phi=\phi_{0}$ in all the region outside the horizon which implies the absence of scalar hair. As a consequence, we find that the $\phi$-dependence in the functions $g_{i}(\phi,X)$ is essential for achieving hairy BH solutions.
@@ IV.1 Hairy BH solution arising from the derivative interaction $\bar{g}_{3}Y$
For the scalar and the axionic-type couplings of the forms $g_{1}F$ and $g_{2}\tilde{F}$, respectively, it has shown that the charged BH solutions with scalar hair exist Gibbons and Maeda (1988); Garfinkle et al. (1991); Lee and Weinberg (1991); Boskovic et al. (2019); Astefanesei et al. (2019); Fernandes et al. (2019). This subsection aims to examine whether the new BH solutions arise from the derivative interaction, $\bar{g}_{3}Y$, in intermediate scales. For this purpose, we concentrate on the case $c_{3}=1$ with vanishing $c_{1}$, $c_{2}$, and $d_{i}$. We solve the field equations given in Eqs. (2.9)-(2.12) numerically with the model parameters $\mu=0.1$ and $P=0.1M_{\rm pl}r_{h}$. The analytic solutions around the horizon given in Eqs. (3.3)-(3.8) are used as boundary conditions with $\phi_{0}=-1.0\times 10^{-4}M_{\rm pl}$ at $r=1.001r_{h}$  [^3] The boundary conditions and model parameters are chosen so that (i) a condition for the couplings to work as corrections to the RN solution, (ii) a condition for avoiding discontinuity, (iii) a condition for the existence of horizon, are satisfied. The concrete expressions of these conditions will be discussed in Sec. IV.2.. We present the result in Fig. 1.
Figure 1: Numerical solutions of $f,h,V,\phi,h-f$ outside the horizon for the coupling $c_{3}=1$, with vanishing $c_{1}$, $c_{2}$, and $d_{i}$. The boundary conditions are chosen to satisfy Eqs. (3.3)-(3.8) with $\mu=0.1$, $P=0.1M_{\rm pl}r_{h}$, $\phi_{0}=-1.0\times 10^{-4}M_{\rm pl}$ at $r=1.001r_{h}$.